# LeanVec: Search your vectors faster by making them fit

### v1

@@ 3       $\gamma\leftarrow 1/(t+1)^{\alpha}$;
6
Algorithm 1 Frank-Wolfe BCD optimization for LeanVec-OOD, Equation 9, with factor $\alpha\in(0,1)$.
In practice, we use early termination in LABEL:algo:leanvec, i.e., we stop the iterations whenever $\left|f\left({\bm{\mathbf{A}}}^{(t+1)},{\bm{\mathbf{B}}}^{(t+1)}\right)-f\left({\bm{\mathbf{A}}}^{(t)},{\bm{\mathbf{B}}}^{(t)}\right)\right|/f\left({\bm{\mathbf{A}}}^{(t)},{\bm{\mathbf{B}}}^{(t)}\right)\leq 10^{-3}$.
open-images-512-1M laion-512-1M
Refer to caption Refer to caption
Figure 2: The Frank-Wolfe algorithm to optimize LeanVec-OOD converges in less than a hundred iterations. In both cases $D=512$. We use $d=128$ for open-images-512-1M (left) and $d=256$ for laion-512-1M (right) and the total runtime is 4 and 13 seconds, respectively. Relaxing the orthogonality constraint is relatively small incurs a relatively small error of $10^{-3}$.
Efficiency. We can precompute ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ at the beginning of the optimization in LABEL:algo:leanvec. This removes the dependency in the number of database and query vectors throughout the iterations and enables dealing with large data and query sets with great efficiency. Moreover, the error between a sample covariance matrix and its expectation converges very quickly (Koltchinskii & Lounici, 2017) with a growing sample size. We can thus safely use a uniform subsampling to compute ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$.
2.4 Convergence analysis of a Frank-Wolfe algorithm for LeanVec-OOD
We consider the general constrained problem
$\min_{z\in{\mathcal{D}}}f(z),$ (15)
We now present two extensions of $\|\nabla f(z^{(t)})\|$ and the Lipschitz assumption, which are standard in unconstrained optimization, suitable for constrained optimization (Lacoste-Julien, 2016).
Definition 1.
The Frank-Wolfe gap at $z^{(t)}$ is defined as
Definition 2.
Lemma 1 (Jaggi (2013, Lemma 7)).
@@ 2.4 Convergence analysis of a Frank-Wolfe algorithm for LeanVec-OOD
In our case, $z=\{{\bm{\mathbf{A}}},{\bm{\mathbf{B}}}\}$, $f$ is defined in Equation 10, and ${\mathcal{D}}=\{z\,|\,{\bm{\mathbf{A}}}\in{\mathcal{C}}\land{\bm{\mathbf{B}}}\in{\mathcal{C}}\}$ for ${\mathcal{C}}$ defined in Equation 3. ${\mathcal{D}}$ is convex, being the intersection of two convex sets, and $f$ is a non-convex function of $z$. The partial derivatives in equations 13 and 14 are linear, and thus L-Lipschitz continuous, and $\operatorname{diam}_{\|\cdot\|}({\mathcal{C}})$ is bounded. This implies that the curvature for ${\bm{\mathbf{A}}}$ (resp. ${\bm{\mathbf{B}}}$), given a fixed ${\bm{\mathbf{B}}}$ (resp. ${\bm{\mathbf{A}}}$) is finite.
We are now ready to state our main convergence result. The proof is given in Appendix B.
Theorem 1.
Interestingly, our setting is very related to one in the recent work by Peng & Vidal (2023). They study Equation 15 with a BDC algorithm, but concentrate on the exact minimization of the subproblems or in a case where retractions are applied. We leave as future work the extension of Theorem 1 and its proof to the general setting with more than two blocks.
@@ 3 Experimental results
${}^{\star}$wit-512-1M $512$ 1M Inner prod. $256$
@@
${}^{\star}$rqa-768-1M $768$ 1M Inner prod. $160$
${}^{\star}$rqa-768-10M $768$ 10M Inner prod. $160$
Datasets. We evaluate the effectiveness of our method on a wide range of datasets with varied sizes ($n=1\text{M}$ to $n=13\text{M}$) and medium to high dimensionalities ($D=200$ to $D=960$), containing in-distribution (ID) and out-of-distribution (OOD) queries, see LABEL:table:datasets. For ID and OOD evaluations, we use standard and recently introduced datasets (Zhang et al., 2022; Babenko & Lempitsky, 2021; Schuhmann et al., 2021; Aguerrebere et al., 2024). We also introduce new datasets with different types of OOD characteristics: cross-modality with wit-512-1M and question-answering with rqa-768-1M and 10M. See Appendix D for more details.
Setup. Throughout the experiments, LeanVec uses LVQ8 for the primary vectors and FP16 for the secondary vectors. For each dataset, we use the dimensionality $d$ that yields the highest search performance at 90% accuracy (see LABEL:table:datasets). To prevent overfitting, we use two separate query sets (see Appendix D): one to learn the LeanVec-OOD projection matrices and to calibrate the runtime search parameters in SVS, and one to generate our results. As standard (Aumüller et al., 2020a), we report the best out of 10 runs for each method. Further experimental details can be found in Appendix C.
Search performance. For this study, we use the graph built with uncompressed FP16 vectors to evaluate the search performance gains provided by LeanVec over the state-of-the-art methods in SVS on high-dimensional embedding vectors.  Figure 3 and  Figure 4 show the search performance on datasets with in-distribution (ID) and out-of-distribution queries, respectively. In the ID datasets, both LeanVec-ID and LeanVec-OOD show similar performance, confirming Proposition 1 in practice. LeanVec-OOD provides up to 10.2x and 3.7x performance gains over FP16 and LVQ, respectively, at a 10-recall@10 of 0.90 on gist-960-1M as it has the highest dimensionality amongst the evaluated datasets. LeanVec-OOD shines on the OOD datasets, outperforming LeanVec-ID and LVQ by up to 1.5x and 2.8x, respectively, at a 10-recall@10 of 0.90 on rqa-768-1M. Note that the performance advantage of LeanVec diminishes when the dimensionality of the dataset is small, as in the case of t2i-200-1M. Lastly, LeanVec does not show significant gains in laion-512-1M. In this case, linear dimensionality reduction significantly impacts the accuracy of the search. We plan to address this issue in future work.
@@
Comparison with the state of the art. In addition to the state-of-the-art SVS-LVQ (Aguerrebere et al., 2023), we compare LeanVec to three widely adopted methods: HNSWlib (Malkov & Yashunin, 2018), Vamana (Jayaram Subramanya et al., 2019), and FAISS-IVFPQfs (Johnson et al., 2021). See Appendix C for further experimental details and configurations. Here, we use LeanVec-OOD as it achieves equal or better performance than LeanVec-ID in all cases. As shown in Figure 6, the combination of LeanVec with the SVS library achieves a significant performance lead over the other prevalent similarity search methods on high-dimensional datasets. SVS-LeanVec provides 1.1x, 1.9x, 2.8x, and 3.7x performance boost on t2i-200-1M, deep-256-1M, rqa-768-1M, and gist-960-1M, respectively, at a 10-recall@10 of 0.90 over the second-best method, SVS-LVQ, and 2.4x, 3.8x, 7.8x, and 8.5x, respectively, over the third-best method, FAISS-IVFPQfs. Note that the advantage gets higher as the dimensionality increases.
@@
LeanVec scaling on larger datasets. We run LeanVec on two larger datasets of 13 and 10 million vectors, namely, open-images-512-13M and rqa-768-10M. For these larger datasets, LeanVec-OOD uses projection matrices of their 1M versions (open-images-512-1M and rqa-768-10M, respectively). As shown in Figure 7, LeanVec continues to show performance gains in larger datasets. LeanVec-OOD achieves 2x and 2.4x performance benefits over LVQ in open-images-512-13M and rqa-768-10M, respectively, at a 10-recall@10 of 0.90. Leaving SVS-LVQ aside, LeanVec-OOD provides much higher benefits when compared to the next best performing methods: 7.9x and 13.7x over HNSWlib in open-images-512-13M and rqa-768-10M, respectively.
Figure 7: LeanVec continues to exhibit superior performance on larger datasets. SVS-LeanVec-OOD achieves performance gains of up to 2.4x over SVS-LVQ and 13.7x over HNSWlib at a 10-recall@10 of 0.90.
@@
Figure 8: The level of dimensionality reduction at which LeanVec provides the best performance varies for each dataset. There is a tradeoff between the accuracy and search performance at different target dimensionalities $d$. In most cases, setting $d=256$ is a reasonable compromise with 2-3x bandwidth reduction.
@@ 4 Related Work
The application of linear dimensionality reduction for approximate nearest neighbor search is not new (Deerwester et al., 1990; Ailon & Chazelle, 2009). However, there are no systematic studies of its effects when applied to deep learning embedding vectors in modern graph-based indices.
@@
Product Quantization (PQ) (Jégou et al., 2011) and other related methods (Ge et al., 2013; Babenko & Lempitsky, 2014; Zhang et al., 2014; André et al., 2015; Matsui et al., 2018; Guo et al., 2020; Wang & Deng, 2020; Johnson et al., 2021; Andre et al., 2021; Ko et al., 2021) were introduced to handle large datasets in settings with limited memory capacity (e.g., Jayaram Subramanya et al., 2019; Jaiswal et al., 2022). However, when used for high-throughput graph-search, with its predominant random memory access pattern, these quantizers do not enable extremely fast similarity computations.
Dimensionality reduction is deeply related to metric learning Bellet et al. (2013). In the ID case, any metric learned for the main dataset will be equally suitable for similarity search. However, this metric may be unsuitable for similarity search in the OOD case. As an instance of deep metric learning (Kaya & Bilge, 2019), Zhang et al. (2022) proposed CCST that uses transformers to reduce the dimensionality of deep learning embedding vectors. However, the computational complexity of transformers precludes their usage for search and circumscribes their application to index construction, where they lead to significant performance gains. As we show in the experimental section, LeanVec outperforms CCST for index construction and can be equally used for search.
Lastly, He et al. (2021) and Izacard et al. (2020) used principal components analysis (PCA) in the context of retrieval-augmented language models (LM), showing that the LM maintains its perplexity score and its overall speed is boosted. However, they treat the similarity search system as a black box and do not address out-of-distribution aspects.

### v2

@@ 4       $\gamma\leftarrow 1/(t+1)^{\alpha}$;
8
Algorithm 1 Frank-Wolfe BCD optimization for Equation 9 with factor $\alpha\in(0,1)$.
Figure 2: LABEL:algo:leanvec converges in 51 iterations for open-images-512-1M with $D=512$ and $d=128$. The total runtime is 4 seconds, respectively. Relaxing the orthogonality constraint incurs a relatively small error of $10^{-3}$.
@@
Equipped with these tools, we can pose the complete optimization procedure in LABEL:algo:leanvec. There, we update ${\bm{\mathbf{A}}}$ (resp. ${\bm{\mathbf{B}}}$) given a fixed ${\bm{\mathbf{B}}}$ (resp. ${\bm{\mathbf{A}}}$) by running one Frank-Wolfe update. The factor $\alpha\in(0,1)$, proposed by Wai et al. (2017) for the step size $\gamma=1/(t+1)^{\alpha}$, can be replaced by a line search to speed up the optimization. In our experiments we did not observe a need for such a performance tuning. In practice, we use early termination in LABEL:algo:leanvec, i.e., we stop the iterations whenever $\left|f\left({\bm{\mathbf{A}}}^{(t+1)},{\bm{\mathbf{B}}}^{(t+1)}\right)-f\left({\bm{\mathbf{A}}}^{(t)},{\bm{\mathbf{B}}}^{(t)}\right)\right|/f\left({\bm{\mathbf{A}}}^{(t)},{\bm{\mathbf{B}}}^{(t)}\right)\leq 10^{-3}$, yielding a fast runtime, see Figure 2. In Appendix C, we prove the convergence rate of LABEL:algo:leanvec to a stationary point of Equation 9.
2.4 Optimizing the LeanVec-OOD loss with eigenvector search
In this section, we assume ${\bm{\mathbf{A}}}={\bm{\mathbf{B}}}$. This assumption leads to a new optimization technique for the LeanVec-OOD loss. Given ${\bm{\mathbf{P}}}={\bm{\mathbf{A}}}={\bm{\mathbf{B}}}$ and eliminating constant factors, Equation 8 can be rewritten as
$\min_{{\bm{\mathbf{P}}}\in\operatorname{St}(D,d)}\operatorname{Tr}\!\left({\bm{\mathbf{P}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{{\bm{\mathbf{P}}}}^{\top}{\bm{\mathbf{P}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}{{\bm{\mathbf{P}}}}^{\top}-2{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}{{\bm{\mathbf{P}}}}^{\top}{\bm{\mathbf{P}}}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}\right).$ (14)
Here, we can see that it would be desirable to align ${\bm{\mathbf{P}}}$ with both the $d$ leading eigenvectors of ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and with those of ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$. An intuitive idea would be to set ${\bm{\mathbf{P}}}$ using the $d$ leading eigenvectors of ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}+{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$.
However, the matrices ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ are summations over two different numbers of samples (i.e., $n$ and $m$ are not necessarily equal). This asymmetry would artificially give more weight, for example, to ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ if $n\gg m$. We compensate this imbalance by scaling the loss Equation 14 by the constant $\frac{1}{nm}$, obtaining
$\min_{{\bm{\mathbf{P}}}\in\operatorname{St}(D,d)}\operatorname{Tr}\!\left({\bm{\mathbf{P}}}\left(\tfrac{1}{m}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}\right){{\bm{\mathbf{P}}}}^{\top}{\bm{\mathbf{P}}}\left(\tfrac{1}{n}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}\right){{\bm{\mathbf{P}}}}^{\top}-2\left(\tfrac{1}{m}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}\right){{\bm{\mathbf{P}}}}^{\top}{\bm{\mathbf{P}}}\left(\tfrac{1}{n}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}\right)\right).$ (15)
Now, we could set ${\bm{\mathbf{P}}}$ to the $d$ leading eigenvectors of $\tfrac{1}{m}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}+\tfrac{1}{n}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$. Although an improvement, this equal weighting is not empirically optimal. We thus add a scalar factor $\beta\in{\mathbb{R}}_{+}$ and examine the eigenvectors of
${\bm{\mathbf{K}}}_{\beta}=\tfrac{1-\beta}{m}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}+\tfrac{\beta}{n}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}.$ (16)
Empirically, we observe in Figure 3 that the loss in Equation 15 is a smooth function of $\beta$ when ${\bm{\mathbf{P}}}\in{\mathbb{R}}^{d\times D}$ is formed by the $d$ leading eigenvectors of ${\bm{\mathbf{K}}}_{\beta}$. Moreover, it has a unique minimizer. Our resulting optimization, summarized in LABEL:algo:leanvec_eig_search, uses a derivative-free scalar minimization technique (Brent, 2013) to find the value of $\beta$ that provides the optimum balance.
1
2algorithm]algo:leanvec_eig_search
3
4Find $\beta\in[0,1]$ that minimizes the loss in Equation 14 with ${\bm{\mathbf{P}}}\leftarrow$ projection($\beta$);
5
6Procedure projection($\beta$)
7       return the matrix ${\bm{\mathbf{P}}}\in\operatorname{St}(D,d)$ formed by the $d$ eigenvectors of ${\bm{\mathbf{K}}}_{\beta}=\frac{1-\beta}{m}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}+\frac{\beta}{n}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ with the largest eigenvalues, where ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}={\bm{\mathbf{Q}}}{{\bm{\mathbf{Q}}}}^{\top}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}={\bm{\mathbf{X}}}{{\bm{\mathbf{X}}}}^{\top}$;
8
Algorithm 2 Eigenvector search optimization for Equation 15.
Figure 3: The loss in Equation 14 is a smooth function of $\beta$ when ${\bm{\mathbf{P}}}=\operatorname{eigsearch}(\beta)$ and has a unique minimizer (different for each $d$). LABEL:algo:leanvec_eig_search finds the minimum (marked with a circle) of this loss. Additional results in Figure 17 of the appendix.
In the ID case, we have $\frac{1}{m}{\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}=\frac{1}{n}{\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ in expectation. The eigenvectors of ${\bm{\mathbf{K}}}_{\beta}$ are invariant to the value of $\beta$. Hence, in this case, LABEL:algo:leanvec_eig_search offers a seamless fallback, becoming equivalent to Equation 4.
LABEL:algo:leanvec_eig_search is highly efficient and achieves good local minima of the LeanVec-OOD loss but, so far, lacks theoretical guarantees. However, we can use LABEL:algo:leanvec to shed light on the quality of the solution ${\bm{\mathbf{P}}}$ given by LABEL:algo:leanvec_eig_search. If we set ${\bm{\mathbf{A}}}^{(0)}\leftarrow{\bm{\mathbf{P}}}$ and ${\bm{\mathbf{B}}}^{(0)}\leftarrow{\bm{\mathbf{P}}}$ in LABEL:algo:leanvec, LABEL:algo:leanvec converges in a handful of iterations, improving the loss by less than 2% as observed in Figure 18 of the appendix (we use line search for the gradient step $\gamma$, to ensure that we stay within the same basin). We observe empirically that the theoretical guarantees of LABEL:algo:leanvec translate to the solutions of LABEL:algo:leanvec_eig_search and posit that further theoretical analysis may help clarify its empirical performance. Lastly, we point out that both algorithms perform similarly in the end-to-end similarity search evaluation (see Figure 18 of the appendix).
@@ 3 Experimental results
t2i-200-10M $200$ 10M Inner prod. $192$
wit-512-1M $512$ 1M Inner prod. $256$
@@
rqa-768-1M $768$ 1M Inner prod. $160$
rqa-768-10M $768$ 10M Inner prod. $160$
Datasets. We evaluate the effectiveness of our method on a wide range of datasets with varied sizes ($n=1\text{M}$ to $n=13\text{M}$) and medium to high dimensionalities ($D=200$ to $D=960$), containing in-distribution (ID) and out-of-distribution (OOD) queries, see LABEL:table:datasets. For ID and OOD evaluations, we use standard and recently introduced datasets (Zhang et al., 2022; Babenko & Lempitsky, 2021; Schuhmann et al., 2021; Aguerrebere et al., 2024). We also introduce new datasets with different types of OOD characteristics: cross-modality with wit-512-1M and question-answering with rqa-768-1M and 10M. See Appendix E for more details.
Setup. Throughout the experiments, LeanVec uses LVQ8 for the primary vectors and FP16 for the secondary vectors. For each dataset, we use the dimensionality $d$ that yields the highest search performance at 90% accuracy (see LABEL:table:datasets). For LeanVec-OOD, we present the results using LABEL:algo:leanvec (LABEL:algo:leanvec_eig_search performs similarly as shown in Figure 13 of the appendix.) To prevent overfitting, we use two separate query sets (see Appendix E): one to learn the LeanVec-OOD projection matrices and to calibrate the runtime search parameters in SVS, and one to generate our results. As standard (Aumüller et al., 2020a), we report the best out of 10 runs for each method. Further experimental details can be found in Appendix D, including a discussion on hyperparameter selection in Section D.2.
Search performance. For this study, we use the graph built with uncompressed FP16 vectors to evaluate the search performance gains provided by LeanVec over the state-of-the-art methods in SVS on high-dimensional embedding vectors. Figures 4 and 5 show the search performance on datasets with in-distribution (ID) and out-of-distribution queries, respectively. In the ID datasets, both LeanVec-ID and LeanVec-OOD show similar performance, confirming Proposition 1 in practice. LeanVec-OOD provides up to 10.2x and 3.7x performance gains over FP16 and LVQ, respectively, at a 10-recall@10 of 0.90 on gist-960-1M as it has the highest dimensionality amongst the evaluated datasets (recall is defined in Section D.3). LeanVec-OOD shines on the OOD datasets, outperforming LeanVec-ID and LVQ by up to 1.5x and 2.8x, respectively, at a 10-recall@10 of 0.90 on rqa-768-1M. Note that the performance advantage of LeanVec diminishes when the dimensionality of the dataset is small, as in the case of t2i-200-1M. Lastly, LeanVec does not show significant gains in laion-512-1M. In this case, linear dimensionality reduction significantly impacts the accuracy of the search. We plan to address this issue in future work.
@@
Comparison with the state of the art. In addition to the state-of-the-art SVS-LVQ (Aguerrebere et al., 2023), we compare LeanVec to three widely adopted methods: HNSWlib (Malkov & Yashunin, 2018), Vamana (Jayaram Subramanya et al., 2019), and FAISS-IVFPQfs (Johnson et al., 2021). See Appendix D for further experimental details and configurations. Here, we use LeanVec-OOD as it achieves equal or better performance than LeanVec-ID in all cases. As shown in Figure 7, the combination of LeanVec with the SVS library achieves a significant performance lead over the other prevalent similarity search methods on high-dimensional datasets. SVS-LeanVec provides 1.1x, 1.9x, 2.8x, and 3.7x performance boost on t2i-200-1M, deep-256-1M, rqa-768-1M, and gist-960-1M, respectively, at a 10-recall@10 of 0.90 over the second-best method, SVS-LVQ, and 2.4x, 3.8x, 7.8x, and 8.5x, respectively, over the third-best method, FAISS-IVFPQfs. Note that the advantage gets higher as the dimensionality increases.
@@
LeanVec scaling on larger datasets. We run LeanVec on three datasets of 13 and 10 million vectors: open-images-512-13M, rqa-768-10M, and t2i-200-10M. As shown in Figure 8, LeanVec continues to show performance gains in larger datasets. LeanVec-OOD achieves 2x and 2.4x performance benefits over LVQ in open-images-512-13M and rqa-768-10M, respectively, at a 10-recall@10 of 0.90. Leaving SVS-LVQ aside, LeanVec-OOD provides much higher benefits when compared to the next best performing methods: 7.9x and 13.7x over HNSWlib in open-images-512-13M and rqa-768-10M, respectively. On t2i-200-10M, the benchmark dataset for the OOD track of the NeurIPS’23 Big-ANN competition (Simhadri et al., 2024), we consider the track winner RoarANN (Chen et al., 2024). SVS-LeanVec-OOD and SVS-LVQ perform similarly due to the original low-dimensionality ($D=200$) and outperform RoarANN by 2x at a 10-recall@10 of 0.90.
Figure 8: LeanVec exhibits superior performance on large-scale datasets. SVS-LeanVec-OOD achieves performance gains of up to 2.4x over SVS-LVQ and 13.7x over HNSWlib at a 10-recall@10 of 0.90. On t2i-200-10M, compared to RoarANN (Chen et al., 2024), the OOD track winner of the NeurIPS’23 Big-ANN competition (Simhadri et al., 2024), SVS-LeanVec and SVS-LVQ provide 2x performance gain at a 10-recall@10 of 0.90.
@@
Figure 9: The level of dimensionality reduction at which LeanVec provides the best performance varies for each dataset. However, in all cases, LeanVec brings clear performance improvements over SVS-LVQ, a state-of-the-art solution that does not involve dimensionality reduction. There is a tradeoff between the accuracy and search performance at different target dimensionalities $d$. In most cases, setting $d=256$ is a reasonable compromise with 2-3x bandwidth reduction.
@@ 4 Related Work
The application of linear dimensionality reduction for approximate nearest neighbor search is not new (Deerwester et al., 1990; Ailon & Chazelle, 2009). A few studies (Jegou et al., 2010; Gong et al., 2012; Babenko & Lempitsky, 2014b; Wei et al., 2014) used it for ID queries while the OOD case has been largely ignored.
@@
Product Quantization (PQ) (Jégou et al., 2011) and other related methods (Ge et al., 2013; Babenko & Lempitsky, 2014a; Zhang et al., 2014; André et al., 2015; Matsui et al., 2018; Guo et al., 2020; Wang & Deng, 2020; Johnson et al., 2021; André et al., 2021; Ko et al., 2021) were introduced to handle large datasets in settings with limited memory capacity (e.g., Jayaram Subramanya et al., 2019; Jaiswal et al., 2022). With these techniques, the similarity between (partitions of) the query and each corresponding centroid is precomputed to create a look-up table of partial similarities. The complete similarity computation can then be posed as a set of indexed gather and accumulate operations on this table, which are generally quite slow (Pase & Agelastos, 2019). This is exacerbated with an increased dimensionality $D$: the lookup table does not fit in L1 cache, which slows down the gather operation even further. Quicker ADC (André et al., 2021) offers a clever fix by optimizing these table lookup operations using AVX shuffle and blend instructions to compute the similarity between a query and multiple database elements in parallel. This parallelism can only be achieved if the database elements are stored contiguously in a transposed fashion. This transposition, and Quicker ADC by extension, are ideally suited for inverted indices (Johnson et al., 2021) but are not compatible with the random memory access pattern in graph-based similarity search.
Dimensionality reduction is deeply related to metric learning (Bellet et al., 2013). In the ID case, any metric learned for the main dataset will be equally suitable for similarity search. However, this metric may be unsuitable for similarity search in the OOD case. As an instance of deep metric learning (Kaya & Bilge, 2019), CCST (Zhang et al., 2022) uses transformers to reduce the dimensionality of deep learning embedding vectors. However, the computational complexity of transformers precludes their usage for search and circumscribes their application to index construction, where they lead to significant performance gains. LeanVec outperforms CCST for index construction (Figure 11) and can be equally used for search.
Lastly, He et al. (2021) and Izacard et al. (2020) used PCA in the context of retrieval-augmented language models, showing that their perplexity score is maintained and their overall speed is boosted. However, they treat the similarity search system as a black box and do not address out-of-distribution aspects.